# Bimanual Manipulation of Steady Hand Eye Robots with Adaptive Sclera Force Control: Cooperative vs. Teleoperation Strategies

### v1

@@ III-A Experimental Setup
The setup components comprised the SHER 2.1 on the right holding the primary force-sensing tool, the SHER 2.0 on the left holding the secondary force-sensing tool, two robot controllers (Galil motion controllers, CA, USA), two PHANTOM Omni robots (SensAble Technologies Inc., MA, USA), two ATI force sensors (Nano17, ATI Industrial Automation, Apex, NC, USA), two force-sensing tools with three optical fiber channels attached to each tool shaft and a total of nine FBG sensors (Technica Optical Components, China) on the three channels, an FBG interrogator (HYPERION si155, Luna) for the primary force-sensing tool and another FBG interrogator (sm130, Micron Optics, Inc.) for the secondary fore-sensing tool, an eye phantom, an armrest, and a surgical microscope (Leica Microsystems) (see Fig.2). All these devices were interconnected to two main computers for the two robots through a TCP-IP connection.
Two 6-DoF ATI Nano 17 force sensors are integrated into the SHERs’ end-effector to measure the user hand’s force and torque exerted on the robot handle and to control the robot in the cooperative mode with an admittance-based control algorithm [7]. For measuring the sclera force applied to the eye phantom by the two surgical instruments, two FBG force-sensing tools are employed. The tool shaft features three FBG fibers arranged at 120-degree intervals, housing a total of nine FBG sensors distributed along the three fibers [19] (see Fig. 2). Optical signals from these FBG sensors are received by the two FBG interrogators, which then convert them into force readings at the sclera point and the tooltip. The force read at the sclera point has two components, $F_{sx}$ and $F_{sy}$, perpendicular to the tool shaft and are used in the adaptive sclera force control algorithm to always maintain the sclera force level below a safe threshold. The two PHANTOM Omnis serve as a joystick controller for the bimanual teleoperation of SHER 2.0 and SHER 2.1. The eye phantom, designed as an anatomical eye model replicating critical features of a natural human eye, is utilized in this experiment.
@@ III-B Experimental Procedure
To assess the performance of the two bimanual control modes, BMAC and BMAT, the user is asked to conduct a vessel-following experiment in sitting and standing postures for each control mode (see Fig. 3a-d). The task involves tracking a colored trajectory on the retinal surface of an eye phantom that simulates retinal veins. The colored trajectory comprises four different colors: red, green, blue, and yellow. A pin is attached at the end of each colored vessel to specify a target point and the user is asked to touch the pin with the tooltip. This requires pinpoint accuracy in tool use by the user. The user starts the experiment by approaching the tip of the primary tool, manipulated by the dominant hand, to the start point at the intersection of the four colored vessels, tracking the four colored vessels in a random order, and touching the target pins at the end of each colored vessel. Once all colors are tracked, the user returns to the intersection point (see Fig. 3e). This procedure is considered one trial. More specific details of the procedure are outlined in [17]. The experiment includes 5 sets of tests collected on five different days. For each test, 10 trials are repeated following a random order of colored vessels, resulting in 50 trials for each mode, namely, standing BMAC, sitting BMAC, standing BMAT, and sitting BMAT. To optimize visualization, all trials are conducted under a surgical microscope with an illuminator.
In the bimanual cooperative control mode, the user holds the tool handle, and the robots are manipulated in a direct cooperation manner (Figs. 3a, 3c). In contrast, there is an indirect interaction between the user and the SHERs through the PHANTOM Omni interfaces in the bimanual teleoperation mode. Two foot-pedals separately activate each robot. The more the foot pedal is pushed, the less stiffness is observed in the SHER handle and the surgical instrument. During the sitting BMAC and sitting BMAT, the user’s hand is stabilized by an armrest to further reduce the hand tremor (see Figs. 3c, 3d), whereas, this armrest is lacking in the standing postures. Regardless of the control mode, the user uses the dominant hand to manipulate the primary force-sensing tool and the non-dominant hand to manipulate a secondary tool for re-orientation of the eye phantom to reach optimum visualization under the microscope. It’s noteworthy that bimanual manipulation is a standard method of performing retinal surgery. All kinematic and force information measured from the robots, the ATI force sensors, and the FBG-based force-sensing tools for each vessel-following trial are recorded in separate .csv files and subsequently analyzed using MATLAB. The statistical analyses were performed using a paired t-test, MATLAB’s ttest2() function, assuming unequal variance (Welch’s t-test). Statistical significance is inferred when p$<$0.05.
@@
Figure 4: An example of the performance of the adaptive force control algorithm on limiting the sclera force components $F_{sx}$ and $F_{sy}$ and sclera force norm $F_{s}$ for the dominant hand (a) and the non-dominant hand (b) during sitting bimanual adaptive teleoperation control mode. Upon reaching the 100 mN threshold, the adaptive sclera force control algorithm is triggered, initiating efforts to minimize the sclera force components, namely $F_{sx}$ and $F_{sy}$, along a desired exponentially decaying force trajectory (dashed curves). The objective of the adaptive force controller is to keep the norm of sclera force $F_{s}$ below a safe threshold of 120 mN [13]. The four peaks in the tip force for the dominant hand show that the four target points (pins) at the end of the colored vessels are touched by the tooltip.
@@
Figure 5: The mean of the tool-sclera interaction forces was calculated over 50 trials for the sitting BMAT mode for (a) dominant hand and (b) non-dominant hand.
@@
TABLE I: Mean and maximum sclera force, handle force/torque, mean completion time, and the percentage of time the sclera force is over 120 mN. The results are reported for the BMAC and the BMAT control modes, for both sitting and standing postures and for both the dominant hand (right robot, SHER 2.1) and non-dominant hand (left robot, SHER 2.0), during the vessel-following experiment. The values inside the parenthesis are the standard deviations.
@@
p-value
Mean sclera force
Max sclera force
Mean completion time
@@
p-value
Mean sclera force
Max sclera force
Mean completion time
@@
p-value
Mean sclera force
Max sclera force
Mean completion time
@@
DH
0.0012
1.564e-08
0.0063
0.0040
NDH
0.0187
0.1621
0.0063
0.1732
IV Experimental Results and Discussion
To the best of our knowledge, the implementation of a bimanual adaptive teleoperation control mode integrated with an adaptive sclera force control algorithm on a microsurgical robotic system used for retinal microsurgery applications, in specific on the SHER 2.0 and SHER 2.1, is unprecedented. The performance of the proposed BMAT mode is compared with a BMAC mode in sitting and standing postures for the user by considering several important metrics related to patient safety and surgeons’ comfort. For example, some factors, having to do with patient safety, including mean and maximum sclera force and the percentage of aggregate completion times in which the recorded sclera force exceeded the safety threshold of 120 mN are reported for both dominant and non-dominant hand for all four control modes. Other factors related to surgeons’ intuition and comfort are also reported such as human-robot interaction force/torque and task completion time (Table I). The completion time is also of great importance which determines the general operational time and cost. Of note, these factors are analyzed and reported for a single expert user and may not be extrapolated to non-expert users who are not at their learning curve plateau [20].
@@ IV-A Sclera Force Control Performance
The adaptive sclera force control algorithm dynamically adjusts the SHER 2.0 and SHER 2.1 end-effector velocity to maintain the sclera force below a safe threshold of 120 mN. This feature enables the surgeon to safely manipulate the eye without applying unsafe force to the sclera. Using this technique, it is possible to safely perform bimanual robot manipulation of the eye without registering the two robots, yet avoid stretching the eye during bimanual manipulation. Figure 4 describes an example of the functionality of the sclera force control algorithm during sitting BMAT control mode for both dominant and non-dominant hands. It is observed that once the sclera force $F_{s}$ reaches the 100 mN threshold, the adaptive sclera force control algorithm gets activated and tries to minimize the sclera force components, $F_{sx}$ and $F_{sy}$, along a desired exponentially decaying force trajectory. The proposed safe limit for sclera force is 120 mN [13]. The 100 mN threshold is selected for activating the adaptive force control to provide a 20 mN offset before the safe limit. The presence of four peaks in the tip force data for the dominant hand (Fig. 4a) indicates that the tooltip made contact with the four target points (pins) located at the ends of the colored vessels. Also, an example of the mean sclera force for all 50 trials for the sitting BMAT control mode is provided in Fig. 5, which indicates that the user showed a consistent performance in this mode for both dominant and non-dominant hands. Similar performance is observed for the other three control/posture modes.
Here, a comparison is made between the standing and sitting postures for each of the two BMAC and BMAT control modes, and also between the sitting BMAC and the sitting BMAT (p-values are reported in Tables IIIV) to compare the performance of BMAC and BMAT and to examine the effects of user’s posture on their performance.
@@
For example, by comparing the standing BMAC and sitting BMAC, it is observed that the user applied a significantly smaller maximum force of 185.14 $\pm$21.93 by his dominant hand during the sitting BMAC compared to this value in the standing BMAC which is 245.57 $\pm$23.07 (p $\ll$ 0.05). This difference is also significant for the non-dominant hand. The percentage of aggregate time when the sclera force is above 120 mN for the dominant hand is 0.45 $\pm$ 0.07 for the sitting BMAC and 2.14 ±$\pm 0.06$ for the standing BMAC (p $\ll$ 0.05). Task completion time is also significantly reduced from 30.81 $\pm$5.35 for the standing BMAC to 23.21 $\pm$14.8 for the sitting BMAC (p = 0.0084) (Tables I, II).
@@
Similarly, by comparing the sitting BMAT and standing BMAT, it is observed that the user demonstrated better performance sitting. For example, the mean sclera force for the non-dominant hand during sitting BMAT is 44.84 $\pm$5.39, whereas this value for the standing BMAT is 62.77 $\pm$22.26 ($\ll$0.05), or the maximum sclera force for the dominant hand in the standing BMAT is 290.49 $\pm$38.89 while this value is reduced to 189.92 $\pm$17.64 for the sitting BMAT ($\ll$0.05). Similar improvement is also noticed for the mean completion time and the percentage of aggregate time when sclera force is above 120 mN (see Tables I, III). Based on these findings, we conclude that the user outperforms in the sitting posture because the arms are stabilized using an armrest, reducing hand tremors and fatigue. Furthermore, it is difficult to handle both foot pedals in standing mode as the user is supposed to hold their weight on the heel and pivot to push the pedals. Sometimes the user did not handle the pedals efficiently, especially the left one which is related to the non-dominant hand. This resulted in higher stiffness of the robots and, therefore, higher maximum sclera force. Although, the sitting posture allows a superior performance that makes it a preferred option for robot-assisted retinal microsurgery, however, the standing posture also allows acceptable and safe performance, which could potentially be used for other surgeries during which the surgeons may decide to partly stand during the operation such as microvascular Anastomosis [21].
@@
Finally, by comparing the results of the BMAT and the BMAC modes in the sitting posture (Tables I, IV), it is observed that bimanual adaptive teleoperation showed comparable or better performance to the bimanual adaptive cooperative mode. For example, the mean sclera force for the dominant hand for the sitting BMAT is 32.04 $\pm$4.84, which is less than the value for the sitting BMAC (35.79 $\pm$5.84, p = 0.0012). Also, the percentage of aggregate time with a scleral force higher than 120 mN is less for the sitting BMAT (0.01 $\pm$7.93e-4) as compared to the sitting BMAC (0.45 $\pm$0.07, p = 0.004). This demonstrates lower scleral forces with a sitting BMAT. This could be, in part, due to the slower end-effector velocity during the teleoperation mode compared to the cooperative mode, which in turn results in a longer task completion time for the sitting BMAT mode (26.53 $\pm$ 6.47 s) compared to the sitting BMAC mode (23.21 $\pm$ 14.8 s, p = 0.0063). The maximum sclera force for both dominant and non-dominant hands for the sitting BMAT is higher than those of the sitting BMAC. For example, for the dominant hand, the value for the sitting BMAC is 185.14 $\pm$ 21.93 mN while this value for the sitting BMAT is slightly higher, 189.92 $\pm$ 17.64 mN. Based on our understanding, this has to do with the repositioning capability of the teleoperation mode. In the cooperative mode, the user can finish the task by holding the pedals during the entire vessel-following trial, whereas, during the teleoperation mode, the motion scaling capability necessitates several repeated unclutching, repositioning of the Omni handles, and clutching to follow a trajectory. This results in a higher stiffness at the SHERs’ handle and the surgical instrument at the beginning of each clutching, which resulted in a higher maximum sclera force for the BMAT mode; however, it is still on the same level of magnitude as the BMAC mode.

### v2

@@ III Experiments
This experiment aims to evaluate the performance of the proposed BMAT control mode regarding patient safety in terms of minimizing the tool-sclera interaction forces and avoiding the stretch and compression of the sclera in the absence of registration between the two robots. Also, the performance of the BMAT framework is compared with that of a bimanual adaptive cooperative framework in both sitting and standing postures.
Figure 4: The experiment setup includes the SHER 2.1 and the SHER 2.0, two PHANTOM Omni haptic interfaces, two FBG interrogators, a surgical microscope, and an armrest (left); an eye phantom and two ATI force/torque sensors (top right); two FBG-based force-sensing tools attached to the SHERs’ handles (bottom right).
Figure 5: Vessel-following experiment procedure for four combinations of control modes and user posture including (a) standing BMAC, (b) standing BMAT, (c) sitting BMAC, and (d) sitting BMAT. (e) The user activates each robot using a foot pedal and follows a random order of colored vessels by the dominant (right) hand, touches a target pin at the end of each vessel by the tooltip of the dominant hand, and re-orients the eye phantom using the non-dominant hand under a surgical microscope.
@@ III-A Experimental Setup
The setup components comprised the SHER 2.1 on the right holding the primary force-sensing tool, the SHER 2.0 on the left holding the secondary force-sensing tool, two robot motion controllers (Galil motion controllers, CA, USA), and two PHANTOM Omni robots (SensAble Technologies Inc., MA, USA). The two PHANTOM Omnis serve as a haptic interface for the bimanual teleoperation control of SHER 2.0 and SHER 2.1. Two 6-DoF force/torque sensors (Nano17, ATI Industrial Automation, NC, USA) are integrated into the SHERs’ end-effector to measure the user hand’s force and torque exerted on the robot handle and to control the robot in the cooperative mode with an admittance-based control algorithm [7]. Two FBG-based force-sensing surgical instruments (tools) measure the scleral force applied to the eye phantom. The tool shaft has three FBG fibers arranged at 120-degree intervals, housing a total of nine FBG sensors (Technica Optical Components, China) distributed along the three fibers [26] (see Fig. 4). Optical signals from these FBG sensors are received by two FBG interrogators, (HYPERION SI155, Luna) for the primary force-sensing tool and (SM130, Micron Optics, Inc.) for the secondary fore-sensing tool, enabling the estimation of the tip force, sclera force, and insertion depth by measuring the wavelength shifts of the nine FBG sensors attached to each tool. The force read at the sclera point has two components, $F_{sx}$ and $F_{sy}$, perpendicular to the tool shaft and are used in the adaptive sclera force control algorithm to always maintain the sclera force level below a safe threshold (see Fig. 2). These force-sensing surgical instruments are calibrated using a method outlined by [15]. An eye phantom is designed to simulate the retinal veins of the human eye. An armrest supports the user’s hand during the vessel-following experiment under a surgical microscope (Leica Microsystems) (see Fig.4). All these devices are interconnected to the two robot computers through a TCP-IP connection.
@@ III-B Experimental Procedure
To assess the performance of the BMAC and BMAT control modes, a user, who is an expert in bimanual manipulation of the SHERs, is asked to conduct a vessel-following experiment in sitting and standing postures for each control mode (see Fig. 5a-d). The task involves tracking a colored trajectory on the retinal surface of an eye phantom that simulates retinal veins. The colored trajectory comprises four different colors: red, green, blue, and yellow. A pin is attached at the end of each colored vessel to specify a target point, and the user is asked to touch the pins with the tooltip. This requires pinpoint accuracy in tool use by the user. The user starts the experiment by approaching the tip of the primary tool, manipulated by the dominant hand, to the start point at the intersection of the four colored vessels, tracking the four colored vessels in a random order, and touching the target pins at the end of each colored vessel. Once all colors are tracked, the user returns to the intersection point (see Fig. 5e). This procedure is considered one trial. More specific details of the procedure are outlined in [24]. The experiment includes 5 sets of tests collected on five different days. For each test, 10 trials are repeated following a random order of colored vessels, resulting in 50 trials for each mode: standing BMAC, sitting BMAC, standing BMAT, and sitting BMAT. All trials are conducted under a surgical microscope with an illuminator to optimize visualization.
In the bimanual cooperative control mode, the user holds the tool handle and manipulates the eye robots in a hand-over-hand strategy (Figs. 5a, 5c). In contrast, in the bimanual teleoperation mode, the user indirectly interacts with the SHERs using the PHANTOM Omni interfaces. Two foot-pedals activate each robot separately. The more the foot pedal is pushed, the less impedance is observed in the SHER handle and the surgical instrument.
Figure 6: Comparison of the mean sclera force (a), maximum sclera force (b), and percent of time when the sclera force is over 120 mN (c) for the BMAC and BMAT control modes in sitting and standing postures for the dominant hand (DH) and non-dominant hand (NDH).
(a) DH
(b) NDH
@@
Figure 7: An example of the performance of the adaptive force control algorithm on limiting the sclera force components $F_{sx}$ and $F_{sy}$ and sclera force norm $F_{s}$ for the dominant hand (a) and the non-dominant hand (b) during sitting bimanual adaptive teleoperation control mode. Upon reaching the 100 mN threshold, the adaptive sclera force control algorithm is triggered to minimize the sclera force components, namely $F_{sx}$ and $F_{sy}$, along a desired exponentially decaying force trajectory (dashed curves). The objective of the adaptive force controller is to keep the norm of sclera force $F_{s}$ below a safe threshold of 120 mN [13]. The four peaks in the tip force for the dominant hand show that the four target pins at the end of the colored vessels are touched by the tooltip.
During the sitting BMAC and sitting BMAT, the user’s hand is stabilized by an armrest to further reduce the hand tremor (see Figs. 5c, 5d), whereas this armrest is lacking in the standing postures. Regardless of the control mode, the user uses the dominant hand to manipulate the primary force-sensing tool and the non-dominant hand to manipulate a secondary tool to reorient the eye phantom to reach optimum visualization under the microscope. Of note, bimanual manipulation is a standard method of performing retinal surgery.
All kinematic and force information measured from the robots, the ATI force sensors, and the FBG-based force-sensing tools for each vessel-following trial are recorded in separate .csv files and subsequently analyzed using MATLAB. The statistical analyses were performed using a paired t-test, MATLAB’s ttest2() function, assuming unequal variance (Welch’s t-test). Statistical significance is inferred when p$<$0.05.
IV Experimental Results
To our knowledge, implementing a bimanual adaptive teleoperation control mode integrated with an adaptive sclera force control algorithm on a microsurgical robotic system for retinal microsurgery applications is unprecedented. The performance of the proposed BMAT mode is compared with that of a BMAC mode in sitting and standing postures for the user by considering several important metrics related to patient safety and surgeons’ comfort. For example, some factors having to do with patient safety, including the mean and maximum sclera force and the percentage of aggregate completion times during which the recorded sclera force exceeded the safety threshold of 120 mN, are reported for both dominant and non-dominant hand for all four control modes (see Fig. 6). Other factors related to surgeons’ intuition and comfort are also reported such as human-robot interaction force/torque and task completion time (Table I). Completion time is also important, determining the general operational time and cost. These factors are analyzed and reported for a single expert user and may not be extrapolated to non-expert users who are not at their learning curve plateau [27].
@@ IV Experimental Results
Figure 8: The mean of the tool-sclera interaction forces is calculated over 50 trials for the sitting BMAT mode for (a) dominant hand and (b) non-dominant hand. This figure shows that the user is an expert in using the robots in both control modes.
@@
TABLE I: Mean and maximum sclera force, handle force/torque, mean completion time, and the percentage of time when the sclera force is over 120 mN. The results are reported for the BMAC and the BMAT control modes, for both sitting and standing postures, and for both the dominant hand (right robot, SHER 2.1) and non-dominant hand (left robot, SHER 2.0), during the vessel-following experiment.
@@
DH 2.6583e-10 7.2431e-09 0.0267 0
NDH 2.4950e-04 7.8348e-07 0.0267 0
@@ IV-A Sclera Force Control Performance
The adaptive sclera force control algorithm dynamically adjusts the SHER 2.0 and SHER 2.1 end-effector velocity to maintain the sclera force below a safe threshold of 120 mN. This feature lets the surgeon safely manipulate the eye without applying unsafe force to the sclera. Using this technique, it is possible to safely perform bimanual robot manipulation of the eye without registering the two robots yet avoid stretching the eye during bimanual manipulation. Figure 7 describes an example of the functionality of the sclera force control algorithm during the sitting BMAT control mode for both dominant and non-dominant hands. It is observed that once the sclera force $F_{s}$ reaches the 100 mN threshold, the adaptive sclera force control algorithm gets activated and tries to minimize the sclera force components, $F_{sx}$ and $F_{sy}$, along a desired exponentially decaying force trajectory. The proposed safe limit for sclera force is 120 mN [13]. The 100 mN threshold is selected for activating the adaptive force control to provide a 20 mN offset below the safe limit to give the AFC algorithm enough time to reduce the scleral force before it reaches 120 mN. The presence of four peaks in the tip force data for the dominant hand (Fig. 7a) indicates that the tooltip made contact with the four target points (pins) located at the ends of the colored vessels. Also, an example of the mean sclera force distribution for all 50 trials for the sitting BMAT control mode is provided in Fig. 8, which indicates that the user is an expert in robot use and showed a consistent performance in this mode for both dominant and non-dominant hands. Similar performance is observed for the other three control/posture modes.
Here, the standing and sitting postures for each of the two BMAC and BMAT control modes are compared. Moreover, the sitting BMAC and the sitting BMAT are compared as well (p-values are reported in Tables IIIV), to analyze the feasibility of BMAC and BMAT and examine the effects of user’s posture on their performance.
@@
For example, by comparing the standing BMAC and sitting BMAC, it is observed that the user applied a significantly smaller maximum force of 167.94 $\pm$ 15.77 mN by his dominant hand during the sitting BMAC compared to this value in the standing BMAC which is 197.57 $\pm$ 21.93 mN ($p$ $<$ 0.01) (see Fig. 6b and Tables I, II). This difference is also significant for the non-dominant hand. The percentage of aggregate time when the sclera force is above 120 mN for the dominant hand is 0.26 $\pm$ 0.01 for the sitting BMAC and 2.10 $\pm$ 0.46 for the standing BMAC ($p$ $<$ 0.01). There is no significant difference in task completion time between the standing and sitting BMAC ($p$ $>$ 0.05).
@@
Similarly, by comparing the sitting BMAT and standing BMAT, it is observed that the user demonstrated better performance sitting. For example, the mean sclera force for the dominant hand during sitting BMAT is 31.49 $\pm$ 4.90 mN, whereas this value for the standing BMAT is 40.57 $\pm$ 7.82 mN ($p$ $<$ 0.01) (see Fig. 6a and Tables I and III), or the maximum sclera force for the dominant hand in the standing BMAT is 290.49 $\pm$ 38.89 mN while this value is significantly reduced to 189.92 $\pm$ 18.71 mN for the sitting BMAT ($p$ $<$ 0.01). Similar improvement is also observed for the mean completion time and the percentage of aggregate time when sclera force is above 120 mN. Based on these findings, we conclude that the user outperforms in the sitting posture because the arms are stabilized using an armrest, reducing hand tremors and fatigue. Furthermore, it is difficult to handle both foot pedals in standing mode as the user is supposed to hold their weight on their heels and pivot to push each pedal separately. Sometimes, the user did not handle the pedals efficiently, especially the left one, related to the non-dominant hand. This resulted in higher stiffness of the robots and, therefore, higher maximum sclera force for the non-dominant hand. Perhaps, using a single pedal instead of two to activate both robots may resolve this issue in the standing posture. Although the sitting posture allows a superior performance that makes it a preferred option for robot-assisted retinal microsurgery, the standing posture also allows acceptable and safe performance, which could potentially be used for other surgeries during which the surgeons may decide to partly stand during the operation such as microvascular anastomosis [23].
@@
Finally, by comparing the results of the BMAT and the BMAC modes in the sitting posture (Tables I, IV), it is observed that bimanual adaptive teleoperation showed comparable or better performance, in some terms, to the bimanual adaptive cooperative mode. For example, the mean sclera force for the non-dominant hand for the sitting BMAT is 44.29 $\pm$ 5.27 mN, which is less than this value for the sitting BMAC (52.25 $\pm$ 13.51, $p$ $<$ 0.01). Also, the percentage of aggregate time with a scleral force higher than 120 mN is significantly less for both dominant and non-dominant hands for the sitting BMAT than the sitting BMAC ($p$ $<$ 0.01) (see Fig. 6c). This demonstrates lower scleral forces with a sitting BMAT. This could be, in part, due to the slower end-effector velocity during the teleoperation mode compared to the cooperative mode, which in turn results in a longer task completion time for the sitting BMAT mode (27.53 $\pm$ 6.47 s) compared to the sitting BMAC mode (21.55 $\pm$ 17.47 s, $p$ $<$ 0.05). The maximum sclera force for the non-dominant hand for the sitting BMAT is less than that of the sitting BMAC (165.76 $\pm$ 15.15 mN vs. 184.09 $\pm$ 19.21 mN, $p$ $<$ 0.01), whereas this is the other way around for the dominant hand (189.92 $\pm$ 18.71 mN vs. 167.94 $\pm$ 15.77 mN, $p$ $<$ 0.01). Based on our understanding, this has to do with the repositioning capability of the teleoperation mode. In the cooperative mode, the user can finish the task by holding the pedals during the entire vessel-following trial, whereas, during the teleoperation mode, the motion scaling capability necessitates, mainly for the dominant hand, several repeated unclutching, repositioning of the Omni handles, and clutching again to follow a trajectory. This results in a higher stiffness at the SHERs’ handle and the surgical instrument at the beginning of each clutching, which resulted in a higher mean and maximum sclera force for the dominant hand for the BMAT mode; however, it is still in a safe range and on the same level of magnitude as the BMAC mode.
V Discussion and Future Work
We proposed a bimanual adaptive teleoperation framework equipped with adaptive force control capability using FBG-based force-sensing instruments in this work.
As opposed to other bimanual teleoperation frameworks that typically depend on the robot registration to satisfy the RCM constraints [19], our proposed adaptive teleoperation framework is capable of performing a safe bimanual manipulation of the human eye despite the absence of registration of the two robots. Both surgical tools, including the secondary tool used for rotating the eyeball, are force-sensing tools that provide sclera force measurements for the adaptive force control algorithm to adjust scleral forces automatically to maintain them below a safe threshold.